### v1

# Establishing a Connection Between the Jet and the Corona in Black Hole Low Mass X-Ray Binaries

Abstract
Low-mass black hole X-ray Binaries (LMXBs) undergo outbursts, during which their brightness increases greatly for timescales of months. The X-ray accretion and radio jet properties change dramatically throughout an outburst in a broadly consistent way between sources. Changes to the accretion flow and the corona are evident through X-ray spectral variations, while the jet’s evolution produces changes in the radio. Typically, high energy emission from the corona initially dominate the X-ray spectrum, and quasi-steady compact jets are observed in the radio. As the outburst progresses, emission from the corona fades and is superseded by lower energy X-ray accretion disk emission. During this transition, the compact jets are quenched and discrete ejecta, called transient jets, are launched. The concurrence of the corona’s weakening and the jet’s transition from compact to transient implies a connection, but the precise relationship has not been established. Motivated by this, we aim to investigate the corona-jet connection. We perform spectral modeling in the hard and soft X-ray, utilizing NuSTAR, NICER, and Swift/XRT observations to track the evolving X-ray corona for three LMXBs: MAXI J1348-630, MAXI J1535-571 and MAXI J1820+070. We use prior work to mark the presence of compact jets and the dates of discrete jet ejections. We find a clear connection between the evolution of the corona and the jet: across all three sources an increase in the distance of the corona from the black hole occurs near the time that the compact jet is quenched and the transient jet is launched.
1 Introduction
Black Hole Low-Mass X-ray Binaries (LMXBs) play an important role in understanding the universal phenomenon of accretion/ejection. During outburst cycles, LMXBs evolve rapidly compared to their more massive analogs (e.g. supermassive black holes) (e.g., Tetarenko et al., 2016; Corbel et al., 2013), making them ideal laboratories to study the accretion/ejection phenomena.
As their outburst cycle progresses, LMXBs show characteristic X-ray spectral states, namely hard (HS), soft (SS), and intermediate states (IS) (e.g., Remillard & McClintock, 2006; Belloni, 2010). LMXB outbursts typically begin in the HS and transition into the SS via the IS before evolving back into the HS through a second IS. These states are thought to be connected to the geometry of the accretion flow near the black hole (e.g., Meyer-Hofmeister et al., 2009). During a typical outburst, LMXBs launch two types of jets: compact jets and transient jets, which are distinguished based on their morphology and radio spectrum. During the HS, partially self-absorbed compact jets produce a flat-to-inverted radio to mm spectrum (Hjellming & Johnston, 1988; Harmon et al., 1995; Fender, 2001, 2006; Gallo et al., 2004; Russell et al., 2013, 2014; Tetarenko et al., 2015), such that the spectral index, , is positive, if , where is flux density and is frequency. Near the X-ray peak of the outburst, as the source transitions from HS to the SS via the IS, the compact jet emission is observed to quench (Fender et al., 2004a; Russell et al., 2020), giving way to bright radio flaring, corresponding to the launching of the transient jet (Hjellming & Rupen, 1995; Fender et al., 1999; Fender, 2001; Carotenuto et al., 2021). The transient jets are composed of discrete knots of synchrotron emitting plasma that are launched out along the jet axis at close to the speed of light (Mirabel & Rodríguez, 1994). The transient jets exhibit an optically thin radio spectrum, where (Fender, 2006). Radio flaring is thought to arise as the ejecta interacts with the pre-existing compact jet or the surroundings (Fender, 2006).
In the X-ray band, during the HS and IS, we generally observe back-scattered X-ray emission from the accretion disk, the so-called relativistic X-ray reflection (e.g., Lightman & White, 1988; Ross & Fabian, 1993; Bharali et al., 2019). The illuminating source creating the X-ray reflection is commonly referred to as the “corona”, which may be located in the inner part of the disk close to the black hole. The corona is believed to be formed by magnetically heated electrons, with a mean energy of 50 – 100 keV (e.g., Haardt & Maraschi, 1991, 1993). The accretion disk photons gain energy from the corona via inverse Compton scattering (Haardt & Maraschi, 1991, 1993). Despite significant efforts, the location and true nature of the corona remains ambiguous. Constraints have been placed on the coronal size, mainly from X-ray timing observations, finding the corona to be very compact and located close to the black hole (Kara et al., 2016).
Early work linking the corona and jet by Henri & Pelletier (1991) suggested a two-phase model in which material ejected from the accretion disk heats the corona and accelerates it to relativistic speeds, producing the jets observed from AGN. Subsequent work on dynamic LMXB systems has continued to link the corona and jet. Vadawale et al. (2003) proposed that the ejection of coronal material may give rise to the transient jets. Markoff et al. (2005) found that the jets could produce some of the emissions attributed to the corona. Koljonen et al. (2015) observed a correlation between the X-ray photon index and properties of the radio jet, suggesting a causal connection between the corona and jet. From X-ray reflection fraction analysis, You et al. (2021) postulated that the corona is a standing shock with material flowing through it. Through analysis of the jet and X-ray spectra of three LMXBs, Cao et al. (2022) linked the softening of the X-ray spectrum to changes in the corona-jet system. Zhang et al. (2023) analyzed the energy and power density spectra of the BH-LMXB MAXI J1535–571, finding a vertically extended jet-like corona that contracts before the disappearance of the jet and begins to expand in its absence.
The process by which particles are ejected via jets in LMXBs is not fully understood, complicating our study of the jet-corona relationship. In the case of compact jets, particles are accelerated into a power-law energy distribution through mechanisms such as shocks or magnetic reconnection (Markoff et al., 2001a; E. M. de Gouveia Dal Pino & Lazarian, A., 2005; de Gouveia Dal Pino, E. M. et al., 2010). This typically occurs in the jet launching region or at internal shock fronts along the jet (Tetarenko et al., 2019; Fender et al., 2004b; Malzac, 2014, 2012). Vadawale et al. (2003) proposed that an ejection of the corona during the transition to the SS could interact and shock the compact jet and lead to the creation of transient jets. However, the onset of compact jet quenching has been observed during hard-only outbursts, which fail to enter the soft state and a transient jet was not launched (e.g., Russell et al., 2014, 2015). Nevertheless, it remains possible that the compact jet is quenched by the transient jets, which are created by the ejected corona (Russell et al., 2020).
In the study presented here, we attempt to better understand the relationship between the corona and the jet during a BH-XRB outburst by comparing their evolution. We perform detailed X-ray spectral modeling to find how the extent of the X-ray corona evolves throughout three LMXB outbursts: MAXI J1348–630 (J1348), MAXI J1820+070 (J1820), and MAXI J1535–571 (J1535). We make use of the radio spectral index (), combined with previously determined transient jet ejection dates, to determine the onset of the transient jet. Section 2 outlines and describes the observations used and details our spectral modeling and radio analysis. In Section 3, we briefly outline our results. In Section 4.
2 Observations and Methods
The black hole or black hole candidates J1820, J1348, and J1535 are selected for this study because their recent outbursts are well-observed from radio to X-rays with transient jet ejection dates determined.
For our targets in this study, we require both X-ray spectral data and radio observations to track the evolving corona and jet. We focus our efforts on the detailed spectral modeling of the recent X-ray data and rely on published radio data to study the corresponding jet features. As NuSTAR’s 3.0-78.0 keV spectral range provides good coverage of the reprocessed spectral features, archival NuSTAR observations form the basis of our X-ray spectral analysis, with each NuSTAR observation forming an “epoch”. Where possible, we supplement the NuSTAR data with NICER or Swift/XRT data to provide coverage of the soft X-ray energy range and better constrain the disk and reflection features. For epochs with soft X-ray data, we only use the parts of the observations that occur simultaneously. Epoch dates, spectral states, and observations utilized are given in Table LABEL:results_table.
2.1 X-ray Data Reduction
All X-ray data reduction was performed with Heasoft version 6.32.1, with the CALDB accessed remotely from October through December 2023.
The NuSTAR data reduction was performed with NuSTARdas version 2.2.1. Filtered event files were created using the nupipeline task. The source spectra were extracted from the clean event files after defining a circular region of centered on the source position. Far from the source location, a similar circular region was used to generate the background spectra. The nuproducts task was used to extract the science products, including auxiliary response files, response matrix files, and energy spectra. If the observations overlapped with NICER or Swift/XRT data, we ran the nuproducts task with a custom good-time-interval file (GTI). In our analysis, we considered NuSTAR spectral data in the energy range keV.
All NICER data were reduced using NICERDAS version 2023-08-22_V011a and nicaldbver_xti_20221001. Cleaned event files were created from NICER data using . Source spectra and responses were created from the clean event files using the command with the specified overlapping GTIs. The background was estimated using the 111https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl3-spect/ background model described in Remillard et al. (2022). We use the NICER spectral data in the keV range to avoid several well-studied instrumental and interstellar medium features below this range222https://heasarc.gsfc.nasa.gov/docs/nicer/data_analysis/workshops/NICER-CalStatus-Markwardt-2021.pdf.
For the Swift/XRT data analysis, we used xrtpipeline version 0.13.7 to generate clean event files. Further, we employed xselect to trim the event files to times concurrent with NuSTAR observations. Pileup was expected for many of the Swift/XRT spectra due to the brightness of these sources. To account for this, we used the prescription given by Romano et al. (2006) to remove the brightest source pixels.
Additionally, spectra were also filtered only to include grade 0 events. The xrtmkarf task was then used to create the appropriate response files. Background spectra for the Swift/XRT observations were not included in our analysis, as the fields of view of many of these observations did not possess regions free of contamination from the bright sources. We ignore the Swift/XRT spectra greater than 4 keV as NuSTAR has higher quality spectral coverage at energies above 4 keV.
Finally, we used the task ftgrouppha to bin all resulting spectra using a signal-to-noise ratio of 10.
2.1.1 MAXI J1535–571 Excluded X-ray
Data Initial modeling of our MJD 58010.26 observations of J1535 led to large residuals in the NuSTAR and particularly in NICER data. This epoch possesses three overlapping NuSTAR and NICER orbits. After inspecting the light curve of this observation, we found a sudden increase in the count rates for the two telescopes during the second orbit. While this flare may indeed be real, the study of such flares is not the focus of this paper. Thus, for this epoch, we only used data from the first and third overlapping orbits; this eliminated the residuals. Additionally, large residuals were present for the NICER data from MJD 58016.87 to MJD 58034.26 and were not used. Photon pileup could be the reason behind this because pile-up can become an issue in NICER above Crab (Center, 09/25/2024). XSPEC calculations of the NICER 2–10 keV flux reveal values significantly above this threshold for these epochs. Various studies reported that J1535 was extraordinarily bright in the soft X-ray band during these epochs (e.g., Tao et al., 2018).
2.2 X-ray Spectral Modeling
The geometry of the corona–disk is often inferred from the “reflection” spectrum (Reynolds & Nowak, 2003). Assuming a the corona is a compact source located at some height above the black hole (i.e., a lamp-post geometry), changes to this component of the X-ray spectrum can be connected to the corona’s distance (or height) from the black hole (Dauser et al., 2013). To constrain this height, we model the spectra of each outburst with three components. Firstly, a multi-colour disk blackbody component, described by the diskbb model (Mitsuda et al., 1984; Makishima et al., 1986). Secondly, we account for the primary X-ray continuum and relativistic reflection from the disk by including the lamp-post relxilllp model version 2.2 (Dauser & Garcia, 2022). Finally, galactic absorption along the line-of-sight is accounted for by the tbabs model (Wilms et al., 2000a). To account for any uncertainty in the cross-instrument calibration, we also include a constant model in our modeling that we freeze at one for NuSTAR/FPMA. All modeling was performed using XSPEC version 12.13 (Arnaud, 1996) with wilm abundances (Wilms et al., 2000b) and xsect set to vern (Verner et al., 1996).
Multiple epochs are modeled together within XSPEC to constrain the model parameters better. For J1348, we model all epochs together. For J1820, we model our epochs in 3 separate groups, with MJDs 58191.94-58225.30 (initial HS) constituting group 1, MJDs 58242.27-58327.05 (covering the end of the HS, the IS, and the beginning of the SS) comprising group 2, and MJDs 58343.61-58404.95 (final HS) making group 3. For J1535, we split our epochs into two groups. Here, group 1 covers the initial HS and most of the first IS and consists of MJDs 58004.35-58017.21. The remaining observations (MJDS 58023.11-58090.77) cover the transition back to the HS and then to the SS and form group 2.
We consider the black hole spin, the iron abundance, the inclination angle of the disk, and the column density to be “global” parameters, which we leave free but tie together between all epochs modeled together.
In the relxilllp model, we let the model return the primary continuum and the expected reflection spectrum for a lamp post geometry. The disk temperature, disk normalization, corona height (), inner-disc radius, the power law index, the disk ionization, the cutoff energy, and relxilllp normalization were all left free to vary between epochs but were linked between observations within a given epoch.
We estimated the errors of the best-fit parameters using the Markov Chain Monte Carlo (MCMC) simulations. For each group of epochs, we use XSPEC’s built-in Goodman-Weare MCMC algorithm with 270 walkers and a chain length of 6 million elements with the first 2 million elements burned. Errors for each parameter were estimated from the resulting chains by sorting the chain values and taking the central 90%. The numerical results of our X-ray spectral modeling can be found in Table LABEL:obstable.
2.2.1 MAXI J1820+070 Additional Blackbody Component
Our initial model failed to reproduce the observed spectra in several J1820 epochs. Excess residuals between our data and this model were observed in both NuSTAR and NICER spectra below keV our observations on MJDs 58201.53, 58201.86, 58224.96, 58259.33, 58297.20, 58315.02, 58349.82, and 58389.35. Following the analysis of the same data by Fabian et al. (2020), we add a blackbody component to our model for this source. This extra blackbody component likely arises from disk material emitting from or within the innermost stable circular orbit (ISCO), e.g. the plunging region (Mummery et al., 2024; Fabian et al., 2020). While a more robust model accounting for material within the plunging region has been developed by Mummery et al. (2024), a complete analysis with such a model is beyond the scope of this paper.
2.3 Analysis of Jet Features
We use the radio fluxes from prior work for all three sources to track the jet behavior. For J1348 and J1535, we also use the radio flux densities and spectral indices, , presented by Carotenuto et al. (2021) and Russell et al. (2019), respectively. For J1820, radio spectral indices were calculated by fitting a power law to the quasi-simultaneous radio flux values (taken from a variety of radio telescopes in the 0.1365-350 GHz range) reported by (Broderick et al., 2018; Bright et al., 2020; Echiburú-Trujillo et al., 2024; Shaw et al., 2021; Atri et al., 2020)and compiled in Echiburú-Trujillo et al. (2024). Throughout this work, we interpret a shift from optically thick to thin values of as a quenching of the compact jet, as it would be expected as part of a transition to the transient jet. We additionally use prior work to determine the dates of the launching of discrete jet ejecta. Jet launch dates for J1348, J1535, and J1820 are taken from Carotenuto et al. (2021), Russell et al. (2019), and Wood et al. (2021) respectively.
3 Results
Our X-ray spectral modeling aimed to constrain the height of the corona above the accretion disk and examine its correlation with the jet’s behavior. Figures 1, 2, and 3 show the variation of the coronal height and with time. In all three sources, the corona’s height maintains a low, nearly constant value during the HS (see Table LABEL:results_table). However, around the time when the compact jet is quenched, and a transient jet is launched, the corona height increases markedly.
The results for J1820 are visualized in Figure 1. During its long initial HS, the corona height hovers at relatively low values, between and until MJD 58259 when it jumps up to before returning to by MJD 58297. Following this, the inferred height of the corona increases to on MJD 58306. This happens concurrently with the radio spectrum becoming more optically thin as decreases to on MJD 58305, then to on MJD 58311. These changes also coincide with the jet ejection event on MJD 58306 (Wood et al., 2021). The corona height remains high for the following two epochs before moving back toward the black hole as the SS progresses. During the second IS, the compact jet is detected again, and corona height reaches . In the final HS, as recovers, the corona height returns to low values of about .
A similar pattern is observed in J1348 (Figure 2). Our observations begin in the initial HS with the corona height low ( ). The corona height subsequently increases to . on MJD 58521.07 The increase in coronal height again coincides with ’s decrease from low positive to steep negative values and the ejection of the jet (Carotenuto et al., 2021). The following observations see the corona return to lower values. In this source, the corona height is modeled at during MJD 58577.03; this immediately precedes the projected date of a second ejection of plasma blob (Carotenuto et al., 2021). As returns to positive values during the final HS, we measure low coronal heights again.
J1535 (Figure 3) also exhibits a low corona height during the initial HS. The height quickly increases to a maximum height of as decreases to negative values (Russell et al., 2019). This increase in coronal height again occurs around the time of the jet ejection (Chauhan et al., 2021; Russell et al., 2019). The corona height falls to low values again for the remainder of our epochs.
4 Discussion
Our results suggest a causal relationship between the jet and the corona, revealing a change in the corona geometry contemporaneous with a jet ejection event. As outlined in Section 3, the main outcomes of our work are the following:
1) In the initial HSs, we observed that the X-ray corona remains steadily close to the black hole, and the compact radio jet is also observed to remain mostly unchanged.
2) As the LMXB transitions from the HS to the SS, we found that the corona height increases dramatically. At around the same time, the compact jet emission is quenched and jet ejecta are launched from the system. This jet change is determined from spatially resolved ejecta and a transition to optically-thin synchrotron emission, as emitted by the transient jet.
3) In J1820, as the outburst fades and the source returns to a hard state, the compact jet recovers (Echiburú-Trujillo et al., 2024), the corona height once again increases before returning to the its original low levels at later times.
The idea of a connection between the radio jet and the corona is not new; for over two decades it has been postulated that the jet or jet base is connected to the corona and responsible for the relativistic reflection features in X-ray binaries (Markoff et al., 2001b, 2003; Markoff & Nowak, 2004). In a model proposed by Markoff et al. (2001b), it has been argued that inverse Compton and synchrotron emission from the base of the jet could be appreciably contributing to the non-thermal power-law spectra or even produce the detected X-ray reflection (Markoff & Nowak, 2004).
Various X-ray spectral and timing studies have already noted that the corona is dynamic in nature, i.e., the corona geometry changes with the accretion rate change (e.g., Kara et al., 2019; García et al., 2020; Lucchini et al., 2021). However, the exact nature of this evolution remains unclear. Kara et al. (2019) studied reverberation lags in J1820 during its initial HS and found that the corona gradually contracted as the source softened. Lucchini et al. (2021) further strengthened this notion and suggested a link between jet-base and corona. In MAXI J1836–194, they found that harder states () require a more compact corona/jet-base region close to the black hole, while softer states require acceleration to move downstream. Further studies of quasi-periodic oscillations evolution led some to suggest a dual corona model (García et al., 2020), that is sometimes also seen as a vertically extended corona/jet ecosystem that varies in size (Zhang et al., 2023; Sripada & Kandulapati, 2022; Liao et al., 2024; Liu et al., 2022; García et al., 2022). To account for the effects of an extended corona on the observed spectrum, Lucchini et al. (2023) invoked a dual corona model in which a lower component (several ) dominates the features observed in the time-average spectrum, while the second component, a higher corona (hundreds of ) primarily accounts for the time-variable features. Lucchini et al. (2023) tested this model on J1820 (NICER observation on MJD 58301). The authors found a lower corona at height , and a much fainter upper corona at height . That NICER observation occurs between our observations on MJDs 58297 and 58306, where our single-lamppost corona height went from to . Our simpler model is consistent with these results if our single lamp-post corona height can be seen as a weighted average of their two heights. The success of their model in explaining the both the time-variability and the time-averaged X-ray spectrum points to a likely scenario: that while a core of the corona resides somewhat close to the black hole, elements of it stretch out to greater distances. Both their model and ours act as probes of this more extended corona.
During the HS, the corona height always remained at a low value or nearly constant (see Table LABEL:results_table), and the radio spectrum is consistently flat or inverted. This suggests that the geometry of the corona is close to the black hole at this time. Russell et al. (2020) found that the compact particle-acceleration region of J1535 lies at distances of thousands of from the black hole during the HS. Echiburú-Trujillo et al. (2024) found that this region lies even further away during J1820’s HS. This indicates that the corona and the jet particle acceleration region are two separate entities.
However, we observed an increase in the corona height around the time of the transient jet ejection, suggesting a connection between the two. We note that the transient jet is launched around the same time as the compact jet quenching in our three sources; as such, due to the low cadence of our X-ray data, we cannot precisely identify what the immediate connection is. We do see an increase in the coronal height at the same time as the compact jet is re-established in J1820, suggesting that the connection may be with the compact jet. Either way, this work implies that the corona and the outflow are connected through some physical process. Considering our findings and following the models of Markoff et al. (2001b); Markoff & Nowak (2004), we argue that the corona and the jet acceleration region are coupled, with continuous feedback between them (Malzac, 2014). A magnetic field likely originating from the accretion disk, or threaded in the event horizon could be the mechanism behind this coupling. Parts of the corona could be ejected out far beyond the lamp-post height assumptions and possibly interact with the first acceleration region or jet base through the magnetic field lines (Narayan et al., 2003; Igumenshchev, 2008; Tchekhovskoy et al., 2011; Janiuk & James, 2022).
An alternative explanation could be that a change in the accretion rate close to the HS to SS transition (point of transient jet eject) makes fast-moving matter clump into slow-moving plasma, creating a shock in the accretion disk (Fender et al., 2004c; Narayan et al., 2012; Janiuk & James, 2022). This shock changes the accompanying magnetic field configuration (Narayan et al., 2003; Igumenshchev, 2008; Tchekhovskoy et al., 2011; Janiuk & James, 2022), and leads to a change in the corona geometry, possibly from compact to the lamp-post structure, which leads to a change in the coronal height.
This also suggests that the corona might have a more complex geometry, and it could be a combination of both an extended corona and the near-point source. Rapid changes to the coronal geometry as the jet transitions to transient have already been observed in Swift J1727.8–1613 (Podgorný et al., 2024). The authors also reported a return to pre-outburst polarization levels upon re-entering the HS, indicating that the coronal geometry and nature were similar to those before the outburst (Podgorný et al., 2024). Furthermore, a drastic decrease in polarization during the SS has been widely observed (Mastroserio et al., 2024; Svoboda et al., 2024). The polarization findings also lend support to the notion of a non-trivial coronal geometry; a radially extended portion is especially indicated by polarization (Veledina et al., 2023; Krawczynski et al., 2022; Ingram et al., 2024; Steiner et al., 2024).
Our analysis cannot determine whether the jet ejection or the coronal height increase occurred first or whether the corona is entwined with the compact jet or the transient jet. Jet ejection times are difficult to pinpoint (Wood et al., 2021; Russell et al., 2020; Carotenuto et al., 2021). Additionally, our X-ray spectral epochs are too irregular to track the exact onset of coronal height changes, making it challenging to identify the cause of these geometric changes. Higher-cadence simultaneous observations in radio and X-ray bands could clarify whether the coronal height increase precedes jet ejection, providing new insights.
Acknowledgements
This work is partly supported by the NSF REU Program (Award Number 2244344) at Montana State University. This work is additionally supported by the Montana Space Grant Consortium (Grant Number 80NSSC20M0042). JC and RS acknowledge support from the Leverhulme Trust grant RPG-2023-240.

### v2

@@
 Black Hole Low-Mass X-ray Binaries (LMXBs) play an important role in understanding the universal phenomenon of accretion/ejection.
-During outburst cycles, LMXBs evolve rapidly compared to their more massive analogs (e.g.
-supermassive black holes) (e.g., Tetarenko et al., 2016; Corbel et al., 2013), making them ideal laboratories to study the accretion/ejection phenomena.
-As their outburst cycle progresses, LMXBs show characteristic X-ray spectral states, namely hard (HS), soft (SS), and intermediate states (IS) (e.g., Remillard & McClintock, 2006; Belloni, 2010).
-LMXB outbursts typically begin in the HS and transition into the SS via the IS before evolving back into the HS through a second IS.
-These states are thought to be connected to the geometry of the accretion flow near the black hole (e.g., Meyer-Hofmeister et al., 2009).
+During outburst cycles, LMXBs evolve rapidly compared to their more massive analogues, supermassive black holes (e.g., 2004cbhg.symp..169M; hopkins2005), making LMXBs ideal laboratories to study the accretion/ejection phenomena.
+As their outburst cycles progresses, LMXBs show characteristic X-ray spectral states, namely the low-hard state (LHS), high-soft state (HSS), and the Intermediate State (IMS).
+The IMS can be further subdivided into the hard (HIMS) and soft (SIMS) intermediate states based primarily on spectral timing properties (e.g., Remillard2006; Belloni2010).
+LMXB outbursts typically begin in the LHS and transition through the HIMS and SIMS to the HSS before evolving back to the LHS via the IMS.
+These states are thought to be connected to the geometry of the accretion flow near the black hole (e.g., Meyer-Hofmeister2009).
 During a typical outburst, LMXBs launch two types of jets: compact jets and transient jets, which are distinguished based on their morphology and radio spectrum.
-During the HS, partially self-absorbed compact jets produce a flat-to-inverted radio to mm spectrum (Hjellming & Johnston, 1988; Harmon et al., 1995; Fender, 2001, 2006; Gallo et al., 2004; Russell et al., 2013, 2014; Tetarenko et al., 2015), such that the spectral index, , is positive, if , where is flux density and is frequency.
-Near the X-ray peak of the outburst, as the source transitions from HS to the SS via the IS, the compact jet emission is observed to quench (Fender et al., 2004a; Russell et al., 2020), giving way to bright radio flaring, corresponding to the launching of the transient jet (Hjellming & Rupen, 1995; Fender et al., 1999; Fender, 2001; Carotenuto et al., 2021).
-The transient jets are composed of discrete knots of synchrotron emitting plasma that are launched out along the jet axis at close to the speed of light (Mirabel & Rodríguez, 1994).
-The transient jets exhibit an optically thin radio spectrum, where (Fender, 2006).
-Radio flaring is thought to arise as the ejecta interacts with the pre-existing compact jet or the surroundings (Fender, 2006).
-In the X-ray band, during the HS and IS, we generally observe back-scattered X-ray emission from the accretion disk, the so-called relativistic X-ray reflection (e.g., Lightman & White, 1988; Ross & Fabian, 1993; Bharali et al., 2019).
+During the LHS, partially self-absorbed compact jets produce a flat-to-inverted radio to mm spectrum (Hjellming1988; Harmon1995; Fender2001; Fender2006; gallo_2004; Russell_D_2013; Russell_2014_1836; Tetarenk_2016), such that the spectral index, , is positive, if , where is flux density and is frequency.
+Near the X-ray peak of the outburst, as the source transitions from HIMS to the SIMS, the compact jet emission is observed to quench (fend04; Russell_2020), giving way to bright radio flaring, corresponding to the launching of the transient jet (Hjellming95; Fender_1999; Fender2001; car).
+The transient jets are composed of discrete knots of synchrotron emitting plasma that are launched out along the jet axis at close to the speed of light (mirabel_2004).
+The transient jets exhibit an optically thin radio spectrum, where (Fender2006).
+Radio flaring is thought to arise as the ejecta interacts with the pre-existing compact jet or the surroundings (Fender2006).
+In the X-ray band, we generally observe back-scattered X-ray emission from the accretion disk, the so-called relativistic X-ray reflection (e.g., Lightman1988; Ross1993; Bharali2019).
 The illuminating source creating the X-ray reflection is commonly referred to as the “corona”, which may be located in the inner part of the disk close to the black hole.
-The corona is believed to be formed by magnetically heated electrons, with a mean energy of 50 – 100 keV (e.g., Haardt & Maraschi, 1991, 1993).
-The accretion disk photons gain energy from the corona via inverse Compton scattering (Haardt & Maraschi, 1991, 1993).
+The corona is believed to be formed by magnetically heated electrons, with a mean energy of 50 – 100 keV (e.g., haardt91; Haardt1993).
+The accretion disk photons gain energy from the corona via inverse Compton scattering (haardt91; Haardt1993).
 Despite significant efforts, the location and true nature of the corona remains ambiguous.
-Constraints have been placed on the coronal size, mainly from X-ray timing observations, finding the corona to be very compact and located close to the black hole (Kara et al., 2016).
-Early work linking the corona and jet by Henri & Pelletier (1991) suggested a two-phase model in which material ejected from the accretion disk heats the corona and accelerates it to relativistic speeds, producing the jets observed from AGN.
+Constraints have been placed on the coronal size, mainly from X-ray reverberation studies, finding the corona to be compact and located close to the black hole (Kara2016).
+Other studies aimed at constraining the geometry of the corona have modeled the RMS-variability and phase-lags of quasi-periodic oscillations (QPOs) observed in LMXB spectra in addition to the time averaged spectra.
+Using spectral-timing models, Garcia2021 and bellavita2022 found that the corona in MAXI J1348–630 can be best described as consisting of two components one small and one more extended.
+A study of MAXI J1348–630’s QPOs during the rise of the outburst and reflares also suggests a two Component corona, with one being more horizontally extended (along the disk) corona to explain the high fraction emission reflected by the disk (alabarta2025).
+According to rawat2023, a two-coronal model is also preferred in the case of MAXI J1535–571.
+Studying the QPOs during the SIMS, zhang2023 suggests a vertically extended jet-like corona in the SIMS.
+This more complex and extended coronal geometry, is in agreement with recent results from (IXPE) ixpe, which have measured the X-ray polarization of several galactic black holes and thus provided evidence of a corona that is elongated and oriented along the disk (Veledina_2023_ixpe; cyg_ixpe, e.g.).
+A changing coronal geometry has also been suggested, ma2023 studied QPOs in MAXI J1820+020 in its HIMS-SIMS transition, suggesting that the corona expands horizontally along the disk during the outburst rise and then transforms into a jet-like extended shape in the SIMS,accompanied by a radio flare.
+belloni2024 studied QPOs in GRS 1915+105 to find a rapidly varying corona—with its size varying between and in a single 10.5 hour observation for a BH reid2023.
+Early work linking the corona and jet by gilles1991 suggested a two-phase model in which material ejected from the accretion disk heats the corona and accelerates it to relativistic speeds, producing the jets observed from AGN.
 Subsequent work on dynamic LMXB systems has continued to link the corona and jet.
-Vadawale et al.
-(2003) proposed that the ejection of coronal material may give rise to the transient jets.
-Markoff et al.
-(2005) found that the jets could produce some of the emissions attributed to the corona.
-Koljonen et al.
-(2015) observed a correlation between the X-ray photon index and properties of the radio jet, suggesting a causal connection between the corona and jet.
-From X-ray reflection fraction analysis, You et al.
-(2021) postulated that the corona is a standing shock with material flowing through it.
-Through analysis of the jet and X-ray spectra of three LMXBs, Cao et al.
-(2022) linked the softening of the X-ray spectrum to changes in the corona-jet system.
-Zhang et al.
-(2023) analyzed the energy and power density spectra of the BH-LMXB MAXI J1535–571, finding a vertically extended jet-like corona that contracts before the disappearance of the jet and begins to expand in its absence.
+Vadawale2003 proposed that the ejection of coronal material may give rise to the transient jets.
+markoff_2005 found that the jets could produce some of the emissions attributed to the corona.
+koljonen_2015 observed a correlation between the X-ray photon index and properties of the radio jet, suggesting a causal connection between the corona and jet.
+From X-ray reflection fraction analysis, you2021 postulated that the corona is a standing shock with material flowing through it.
+Through analysis of the jet and X-ray spectra of three LMXBs, cao2022 linked the softening of the X-ray spectrum to changes in the corona-jet system.
+A study of radio and X-ray data of GRS 1915+105 covering a 10 year period also revealed a correlation between the radio flux from the jet and the strength of the Fe K line Mendez2022, which is thought to arise from the X-ray reflection of coronal emission.
+Several studies performing specto-temporal analysis in the HIMS and SIMS of MAXI J1535–571 using Insight–HXMT and NICER observations have found a vertically extended jet-like corona that contracts before the disappearance of the jet and begins to expand in its absence (Zhang2022; zhang2023; rawat2023)
 The process by which particles are ejected via jets in LMXBs is not fully understood, complicating our study of the jet-corona relationship.
-In the case of compact jets, particles are accelerated into a power-law energy distribution through mechanisms such as shocks or magnetic reconnection (Markoff et al., 2001a; E.
-M.
-de Gouveia Dal Pino & Lazarian, A., 2005; de Gouveia Dal Pino, E.
-M.
-et al., 2010).
-This typically occurs in the jet launching region or at internal shock fronts along the jet (Tetarenko et al., 2019; Fender et al., 2004b; Malzac, 2014, 2012).
-Vadawale et al.
-(2003) proposed that an ejection of the corona during the transition to the SS could interact and shock the compact jet and lead to the creation of transient jets.
-However, the onset of compact jet quenching has been observed during hard-only outbursts, which fail to enter the soft state and a transient jet was not launched (e.g., Russell et al., 2014, 2015).
-Nevertheless, it remains possible that the compact jet is quenched by the transient jets, which are created by the ejected corona (Russell et al., 2020).
+In the case of compact jets, particles are accelerated into a power-law energy distribution through mechanisms such as shocks or magnetic reconnection (markoff_2001; dalpino_2005; dalpino_2010).
+This typically occurs in the jet launching region or at internal shock fronts along the jet (Tetarenko_2019; Fender2004; Malzac2014; Malzac_2012).
+Vadawale2003 proposed that an ejection of the corona during the transition to the HSS could interact and shock the compact jet and lead to the creation of transient jets.
+However, the onset of compact jet quenching has been observed during hard-only outbursts, which fail to enter the soft state and a transient jet was not launched (e.g., Russell_2014; Russell2015).
+Nevertheless, it remains possible that the compact jet is quenched by the transient jets, which are created by the ejected corona (Russell_2020).
 In the study presented here, we attempt to better understand the relationship between the corona and the jet during a BH-XRB outburst by comparing their evolution.
@@
 The black hole or black hole candidates J1820, J1348, and J1535 are selected for this study because their recent outbursts are well-observed from radio to X-rays with transient jet ejection dates determined.
+We rely on prior work to determine the spectral states and the dates of transition between them.
+For J1820, we adopt the state transitions dates determined by Shidatsu_2019, and identify the HIMS–SIMS transition by the QPO change reported in (Homan2020).
+For J1348, we use the state transitions outlined in zhang, while for J1535 we take spectral states from Tao_2018 and nakahira_2018 compiled by Russell_2019.
 For our targets in this study, we require both X-ray spectral data and radio observations to track the evolving corona and jet.
@@
 For epochs with soft X-ray data, we only use the parts of the observations that occur simultaneously.
-Epoch dates, spectral states, and observations utilized are given in Table LABEL:results_table.
+Epoch dates, spectral states, and observations utilized are given in Table 1.
 2.1 X-ray Data Reduction
@@
 Source spectra and responses were created from the clean event files using the command with the specified overlapping GTIs.
-The background was estimated using the 111https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl3-spect/ background model described in Remillard et al.
-(2022).
+The background was estimated using the 111https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl3-spect/ background model described in Remillard2022.
 We use the NICER spectral data in the keV range to avoid several well-studied instrumental and interstellar medium features below this range222https://heasarc.gsfc.nasa.gov/docs/nicer/data_analysis/workshops/NICER-CalStatus-Markwardt-2021.pdf.
@@
 Pileup was expected for many of the Swift/XRT spectra due to the brightness of these sources.
-To account for this, we used the prescription given by Romano et al.
-(2006) to remove the brightest source pixels.
+To account for this, we used the prescription given by Romano2006 to remove the brightest source pixels.
 Additionally, spectra were also filtered only to include grade 0 events.
@@
 Additionally, large residuals were present for the NICER data from MJD 58016.87 to MJD 58034.26 and were not used.
-Photon pileup could be the reason behind this because pile-up can become an issue in NICER above Crab (Center, 09/25/2024).
+Photon pileup could be the reason behind this because pile-up can become an issue in NICER above Crab (nicer_pileup).
 XSPEC calculations of the NICER 2–10 keV flux reveal values significantly above this threshold for these epochs.
-Various studies reported that J1535 was extraordinarily bright in the soft X-ray band during these epochs (e.g., Tao et al., 2018).
+Various studies reported that J1535 was extraordinarily bright in the soft X-ray band during these epochs (e.g., Tao_2018).
 2.2 X-ray Spectral Modeling
-The geometry of the corona–disk is often inferred from the “reflection” spectrum (Reynolds & Nowak, 2003).
-Assuming a the corona is a compact source located at some height above the black hole (i.e., a lamp-post geometry), changes to this component of the X-ray spectrum can be connected to the corona’s distance (or height) from the black hole (Dauser et al., 2013).
+The geometry of the corona–disk is often inferred from the “reflection” spectrum (Reynolds2003).
+Approximating the corona as a compact source located at some height above the black hole (i.e., a lamp-post geometry), changes to this component of the X-ray spectrum can be connected to the corona’s distance (or height) from the black hole (dauser).
+While the corona is likely more extended, as indicated by recent polarization studies (see discussion), we adopt this simple lamp-post geometry as a probe of the evolution of the Comptonizing region for its ease of use and interpretation.
 To constrain this height, we model the spectra of each outburst with three components.
-Firstly, a multi-colour disk blackbody component, described by the diskbb model (Mitsuda et al., 1984; Makishima et al., 1986).
-Secondly, we account for the primary X-ray continuum and relativistic reflection from the disk by including the lamp-post relxilllp model version 2.2 (Dauser & Garcia, 2022).
-Finally, galactic absorption along the line-of-sight is accounted for by the tbabs model (Wilms et al., 2000a).
+Firstly, a multi-colour disk blackbody component, described by the diskbb model (diskbb; Makishima1986).
+Secondly, we account for the primary X-ray continuum and relativistic reflection from the disk by including the lamp-post relxilllp model version 2.2 (relxill).
+Finally, galactic absorption along the line-of-sight is accounted for by the tbabs model (tbabs).
 To account for any uncertainty in the cross-instrument calibration, we also include a constant model in our modeling that we freeze at one for NuSTAR/FPMA.
-All modeling was performed using XSPEC version 12.13 (Arnaud, 1996) with wilm abundances (Wilms et al., 2000b) and xsect set to vern (Verner et al., 1996).
+All modeling was performed using XSPEC version 12.13 (xspec) with wilm abundances (wilm) and xsect set to vern (verners).
 Multiple epochs are modeled together within XSPEC to constrain the model parameters better.
 For J1348, we model all epochs together.
-For J1820, we model our epochs in 3 separate groups, with MJDs 58191.94-58225.30 (initial HS) constituting group 1, MJDs 58242.27-58327.05 (covering the end of the HS, the IS, and the beginning of the SS) comprising group 2, and MJDs 58343.61-58404.95 (final HS) making group 3.
+Due to computational limitations arising from the large number of epochs in the other two sources, we modeled these epochs in separate groups rather than all together.
+For J1820, we model our epochs in 3 separate groups, with MJDs 58191.94-58225.30 (initial LHS) constituting group 1, MJDs 58242.27-58327.05 (covering the end of the LHS, the HIMS and SIMS, and the beginning of the HSS) comprising group 2, and MJDs 58343.61-58404.95 (final LHS) making group 3.
 For J1535, we split our epochs into two groups.
-Here, group 1 covers the initial HS and most of the first IS and consists of MJDs 58004.35-58017.21.
-The remaining observations (MJDS 58023.11-58090.77) cover the transition back to the HS and then to the SS and form group 2.
-We consider the black hole spin, the iron abundance, the inclination angle of the disk, and the column density to be “global” parameters, which we leave free but tie together between all epochs modeled together.
-In the relxilllp model, we let the model return the primary continuum and the expected reflection spectrum for a lamp post geometry.
+Here, group 1 covers the initial LHS and most of the transition to the HSS and consists of MJDs 58004.35-58017.21.
+The remaining observations (MJDS 58023.11-58090.77) cover the transition back to the LHS and then to the HSS and form group 2.
+We consider the black hole spin, the iron abundance, the inclination angle of the disk, and the column density to be “global” parameters, which we leave free but tie together between all epochs in each group.
+These global parameters were allowed to vary between groups.
+We discuss the impact of this further in section 3.
+In the relxilllp model, we let the model return the primary continuum and the expected reflection spectrum for a lamp post geometry, while diskbb accounts for the temperature and normalization of the accretion disk.
 The disk temperature, disk normalization, corona height (), inner-disc radius, the power law index, the disk ionization, the cutoff energy, and relxilllp normalization were all left free to vary between epochs but were linked between observations within a given epoch.
@@
 Errors for each parameter were estimated from the resulting chains by sorting the chain values and taking the central 90%.
-The numerical results of our X-ray spectral modeling can be found in Table LABEL:obstable.
+The numerical results of our X-ray spectral modeling can be found in Table 1.
 2.2.1 MAXI J1820+070 Additional Blackbody Component
@@
 Excess residuals between our data and this model were observed in both NuSTAR and NICER spectra below keV our observations on MJDs 58201.53, 58201.86, 58224.96, 58259.33, 58297.20, 58315.02, 58349.82, and 58389.35.
-Following the analysis of the same data by Fabian et al.
-(2020), we add a blackbody component to our model for this source.
+Following the analysis of the same data by Fabian_2020, we add a blackbody component to our model for this source.
 This extra blackbody component likely arises from disk material emitting from or within the innermost stable circular orbit (ISCO), e.g.
-the plunging region (Mummery et al., 2024; Fabian et al., 2020).
-While a more robust model accounting for material within the plunging region has been developed by Mummery et al.
-(2024), a complete analysis with such a model is beyond the scope of this paper.
+the plunging region (1820_isco; Fabian_2020).
+While a more robust model accounting for material within the plunging region has been developed by 1820_isco, a complete analysis with such a model is beyond the scope of this paper.
 2.3 Analysis of Jet Features
 We use the radio fluxes from prior work for all three sources to track the jet behavior.
-For J1348 and J1535, we also use the radio flux densities and spectral indices, , presented by Carotenuto et al.
-(2021) and Russell et al.
-(2019), respectively.
-For J1820, radio spectral indices were calculated by fitting a power law to the quasi-simultaneous radio flux values (taken from a variety of radio telescopes in the 0.1365-350 GHz range) reported by (Broderick et al., 2018; Bright et al., 2020; Echiburú-Trujillo et al., 2024; Shaw et al., 2021; Atri et al., 2020)and compiled in Echiburú-Trujillo et al.
-(2024).
+For J1348 and J1535, we also use the radio flux densities and spectral indices, , presented by car and Russell_2019, respectively.
+For J1820, radio spectral indices were calculated by fitting a power law to the quasi-simultaneous radio flux values (taken from a variety of radio telescopes in the 0.1365-350 GHz range) reported by (atel11887; Bright_2020; Echibur_Trujillo_2024; Shaw_2021; atri2020) and compiled in Echibur_Trujillo_2024.
 Throughout this work, we interpret a shift from optically thick to thin values of as a quenching of the compact jet, as it would be expected as part of a transition to the transient jet.
-We additionally use prior work to determine the dates of the launching of discrete jet ejecta.
-Jet launch dates for J1348, J1535, and J1820 are taken from Carotenuto et al.
-(2021), Russell et al.
-(2019), and Wood et al.
-(2021) respectively.
+We note that although the compact jet is quenched during the HSS, steep-spectrum radio emission can linger at the core of the outburst deep into the HSS (e.g., Russell_2019; hughes_2025a).
+For J1535, this core radio emission faded 5 months into the HSS, while in the case of the black hole low-mass X-ray binary Swift J1727.8–1613, the radio emission from the core remained until the transition back to the LHS (hughes2025b).
+This lingering emission is often attributed to transient ejecta that have not traveled far enough to be resolved from the core position.
+To determine the dates of the launching of discrete jet ejecta we rely on values reported in the literature.
+Jet launch dates for J1535, and J1820 are taken from Russell_2019 and Wood_2021 respectively.
+In the case of MAXI J1348-630, there were multiple excursions to the IMS during the HSS (car; zhang).
+One of these excursions launched a second ejection event that became resolved from the compact source.
+car detected both the first and second ejection events as they propagated away from the compact object.
 3 Results
@@
 Figures 1, 2, and 3 show the variation of the coronal height and with time.
-In all three sources, the corona’s height maintains a low, nearly constant value during the HS (see Table LABEL:results_table).
+In all three sources, the corona’s height maintains a low, nearly constant value during the LHS (see Table 1).
 However, around the time when the compact jet is quenched, and a transient jet is launched, the corona height increases markedly.
 The results for J1820 are visualized in Figure 1.
-During its long initial HS, the corona height hovers at relatively low values, between and until MJD 58259 when it jumps up to before returning to by MJD 58297.
+During its long initial LHS, the corona height hovers at relatively low values, between and until MJD 58259 when it jumps up to before returning to by MJD 58297.
 Following this, the inferred height of the corona increases to on MJD 58306.
 This happens concurrently with the radio spectrum becoming more optically thin as decreases to on MJD 58305, then to on MJD 58311.
-These changes also coincide with the jet ejection event on MJD 58306 (Wood et al., 2021).
-The corona height remains high for the following two epochs before moving back toward the black hole as the SS progresses.
-During the second IS, the compact jet is detected again, and corona height reaches .
-In the final HS, as recovers, the corona height returns to low values of about .
+These changes also coincide with the jet ejection event on MJD 58306 (Wood_2021).
+The corona height remains high for the following two epochs before moving back toward the black hole as the HSS progresses.
+During the transition to the second LHS, the compact jet is detected again, and corona height reaches .
+In the final LHS, as recovers, the corona height returns to low values of about .
 A similar pattern is observed in J1348 (Figure 2).
-Our observations begin in the initial HS with the corona height low ( ).
+Our observations begin in the initial LHS with the corona height low ( ).
 The corona height subsequently increases to .
-on MJD 58521.07 The increase in coronal height again coincides with ’s decrease from low positive to steep negative values and the ejection of the jet (Carotenuto et al., 2021).
+on MJD 58521.07 The increase in coronal height again coincides with ’s decrease from low positive to steep negative values and the ejection of the jet (car).
 The following observations see the corona return to lower values.
-In this source, the corona height is modeled at during MJD 58577.03; this immediately precedes the projected date of a second ejection of plasma blob (Carotenuto et al., 2021).
-As returns to positive values during the final HS, we measure low coronal heights again.
-J1535 (Figure 3) also exhibits a low corona height during the initial HS.
-The height quickly increases to a maximum height of as decreases to negative values (Russell et al., 2019).
-This increase in coronal height again occurs around the time of the jet ejection (Chauhan et al., 2021; Russell et al., 2019).
+In this source, the corona height is modeled at during MJD 58577.03; this immediately precedes the projected date of a second ejection of plasma blob (car).
+As returns to positive values during the final LHS, we measure low coronal heights again.
+J1535 (Figure 3) also exhibits a low corona height during the initial LHS.
+The height quickly increases to a maximum height of as decreases to negative values (Russell_2019).
+This increase in coronal height again occurs around the time of the jet ejection (chauhan_2021; Russell_2019).
 The corona height falls to low values again for the remainder of our epochs.
+In J1820 and J1535, the decision to separate our epochs into groups led to a different set of “global” parameters for each group.
+There are notable variations in some of these parameters between different spectral groupings of the same source.
+Most notable of these differences is the gradual change in inclination through the three groups of J1820 from 26∘ in group 1 to 30∘ in group 2 to 47∘ in group 3.
+Fitting was initially performed with the inclination and other global parameters of groups 1 and 3 frozen to best fit values of the those in group 2.
+This resulted in a poorer , but still preserved our reported trend in coronal height changes.
+While significant changes to the inclination of LMXB GRS 1915+105 have been observed to occur within a year (Rodriguez_2025), this is not something generally observed or expected in LMXBs.
+Another global parameter exhibiting changes between groups is the tbabs equivalent hydrogen column density parameter nH, which goes from about 0.4 atoms cm-2 in group 1 to about 0.07 atoms cm-2 in groups 2 and 3 of J1820.
+A version of our J1820 group 1 fits with the nH parameter fixed to 0.07 was performed with results shown in the appendix alongside our reported results.
+This resulted in a poorer , with no significant change to the corona heights.
+We note that the number of epochs using supplementary NICER or Swift/XRT data differs between groups, potentially contributing to differences in some of the global parameters.
+However, this is unlikely to affect the observed trend in corona heights, as illustrated by our attempts to mitigate the global parameter differences.
+Additionally, the major increases and decreases in corona height are seen to occur within our individual spectral groups.
 4 Discussion
-Our results suggest a causal relationship between the jet and the corona, revealing a change in the corona geometry contemporaneous with a jet ejection event.
+Our results suggest a relationship between the jet and the corona, revealing a change in the corona geometry contemporaneous with a jet ejection event.
 As outlined in section 3, the main outcomes of our work are the following:
-1) In the initial HSs, we observed that the X-ray corona remains steadily close to the black hole, and the compact radio jet is also observed to remain mostly unchanged.
-2) As the LMXB transitions from the HS to the SS, we found that the corona height increases dramatically.
-At around the same time, the compact jet emission is quenched and jet ejecta are launched from the system.
-This jet change is determined from spatially resolved ejecta and a transition to optically-thin synchrotron emission, as emitted by the transient jet.
-3) In J1820, as the outburst fades and the source returns to a hard state, the compact jet recovers (Echiburú-Trujillo et al., 2024), the corona height once again increases before returning to the its original low levels at later times.
-The idea of a connection between the radio jet and the corona is not new; for over two decades it has been postulated that the jet or jet base is connected to the corona and responsible for the relativistic reflection features in X-ray binaries (Markoff et al., 2001b, 2003; Markoff & Nowak, 2004).
-In a model proposed by Markoff et al.
-(2001b), it has been argued that inverse Compton and synchrotron emission from the base of the jet could be appreciably contributing to the non-thermal power-law spectra or even produce the detected X-ray reflection (Markoff & Nowak, 2004).
-Various X-ray spectral and timing studies have already noted that the corona is dynamic in nature, i.e., the corona geometry changes with the accretion rate change (e.g., Kara et al., 2019; García et al., 2020; Lucchini et al., 2021).
-However, the exact nature of this evolution remains unclear.
-Kara et al.
-(2019) studied reverberation lags in J1820 during its initial HS and found that the corona gradually contracted as the source softened.
-Lucchini et al.
-(2021) further strengthened this notion and suggested a link between jet-base and corona.
+1) In the initial LHSs, we observed that the X-ray corona remains steadily close to the black hole, and the compact radio jet is also observed to remain mostly unchanged.
+2) As the LMXB transitions from the LHS to the HSS, we find that the corona height increases dramatically.
+Concurrently, the compact jet emission is quenched, and transient jet ejecta are launched from the system.
+This change in jet behavior is identified through spatially resolved ejecta and the emergence of optically thin synchrotron emission, characteristic of the short-lived transient jet before the jet disappears entirely as the source enters the HSS.
+3) In J1820, as the outburst fades and the source returns to a hard state, the compact jet recovers (Echibur_Trujillo_2024), the corona height once again increases before returning to the its original low levels at later times.
+The idea of a connection between the radio jet and the corona is not new; for over two decades it has been postulated that the jet or jet base is connected to the corona and responsible for the relativistic reflection features in X-ray binaries (Markoff2001; Markoff2003; Markoff2004).
+In a model proposed by Markoff2001, it has been argued that inverse Compton and synchrotron emission from the base of the jet could be contributing to the non-thermal power-law spectra or even produce the detected X-ray reflection (Markoff2004).
+Alternatively, it has been proposed that the Compton-scattering of soft photons could occur entirely within the jet during the LHS (Reig2003; Giannios2004; kylafis2008), and that the lingering power-law component seen during the HSS could be produced by a cool disk extending all the way to the ISCO rather than a corona (Reig2001).
+Comptonization by the jet rather than a corona has been somewhat successful in explaining the reflection component (Reig2021), however this model predicts a flatter emissivity than what reflection model fits typically find (uttley2025).
+Wang_2020hh suggested that there are two Comptonization regions, with one being the more compact corona near the black hole suggested by kara19, and the other occurring within the jet.
+Various X-ray spectral and timing studies have already noted that the corona is dynamic in nature, i.e., the corona geometry changes with the accretion rate change (e.g., kara19; Garc_a_2020; Lucchini2021; alabarta2025; rawat2023; ma2023).
+Studies of QPOs in LMXBs have revealed much about the dynamics of the corona.
+Fast variations in the QPO time lags of GRS 1915+105 suggest that the size of the corona can vary by nearly an order of magnitude over the course of a single 10.5 hour observation (belloni2024).
+However, the exact nature of the corona evolution remains unclear.
+kara19 studied reverberation lags in J1820 during its initial LHS and found that the corona gradually contracted as the spectrum softened.
+We find that our results do not align with those reported by kara19.
+This discrepancy may arise because the two studies are probing different regions of the corona.
+While kara19 observed significant variability of the lags, which they interpreted as reverberation lags, the iron K emission line profile—on which our modeling is based—remains relatively stable within NICER’s energy band.
+Lucchini2021 provided further evidence for a dynamic corona and suggested a link to the jet-base.
 In MAXI J1836–194, they found that harder states () require a more compact corona/jet-base region close to the black hole, while softer states require acceleration to move downstream.
-Further studies of quasi-periodic oscillations evolution led some to suggest a dual corona model (García et al., 2020), that is sometimes also seen as a vertically extended corona/jet ecosystem that varies in size (Zhang et al., 2023; Sripada & Kandulapati, 2022; Liao et al., 2024; Liu et al., 2022; García et al., 2022).
-To account for the effects of an extended corona on the observed spectrum, Lucchini et al.
-(2023) invoked a dual corona model in which a lower component (several ) dominates the features observed in the time-average spectrum, while the second component, a higher corona (hundreds of ) primarily accounts for the time-variable features.
-Lucchini et al.
-(2023) tested this model on J1820 (NICER observation on MJD 58301).
+Further studies of quasi-periodic oscillations evolution led some to suggest a dual corona model (Garcia2021), that is sometimes also seen as a vertically extended corona/jet ecosystem that varies in size.
+(zhang2023; sripada2022; liao2024; Liu_2022; Garcia2022; rawat2023; Mendez2022; ma2023).
+For example, zhang2023 found a vertically extended corona varying from 100 to 700 during the HIMS-SIMS transition of J1535 (assuming a BH).
+liao2024 found corona sizes that varied between and in Swift J1727.8–1613 (for a BH of size (sanchez2025)).
+In a NICER spectral timing study of J1820 wang2021 found increasing soft reverberation lags during the hard-to-soft transition, a change they attribute to the corona expanding from to .
+This trend of an expanding corona is consistent with our results from the same time period, and they suggest this expansion could result in the launching of a jet knot.
+Alternatively, these increasing lags could also be understood in the context of an evolving accretion disk atmosphere, although the thermalization timescale is often very small (salveson2022)
+To account for the effects of an extended corona on the observed spectrum, Lucchini2023 utilized a dual corona model, originally proposed by Garcia2021, in which a lower component (several ) dominates the features observed in the time-average spectrum, while the second component, a higher corona (hundreds of ) primarily accounts for the time-variable features.
+Lucchini2023 tested this model on J1820 (NICER observation on MJD 58301).
 The authors found a lower corona at height , and a much fainter upper corona at height .
 That NICER observation occurs between our observations on MJDs 58297 and 58306, where our single-lamppost corona height went from to .
-Our simpler model is consistent with these results if our single lamp-post corona height can be seen as a weighted average of their two heights.
-The success of their model in explaining the both the time-variability and the time-averaged X-ray spectrum points to a likely scenario: that while a core of the corona resides somewhat close to the black hole, elements of it stretch out to greater distances.
+Our simpler model is consistent with these results if our single lamp-post corona height can be seen as an flux-weighted average position of the two coronal positions.
+The success of their model in explaining both the time-variability and the time-averaged X-ray spectrum points to a likely scenario: that while a core of the corona resides somewhat close to the black hole, elements of it stretch out to greater distances.
 Both their model and ours act as probes of this more extended corona.
-During the HS, the corona height always remained at a low value or nearly constant (see Table LABEL:results_table), and the radio spectrum is consistently flat or inverted.
+During the LHS, the corona height always remained at a low value or nearly constant (see Table 1), and the radio spectrum is consistently flat or inverted.
 This suggests that the geometry of the corona is close to the black hole at this time.
-Russell et al.
-(2020) found that the compact particle-acceleration region of J1535 lies at distances of thousands of from the black hole during the HS.
-Echiburú-Trujillo et al.
-(2024) found that this region lies even further away during J1820’s HS.
+Russell_2020 found that the compact particle-acceleration region of J1535 lies at distances of thousands of from the black hole during the LHS.
+Echibur_Trujillo_2024 found that this region lies even further away during J1820’s LHS.
 This indicates that the corona and the jet particle acceleration region are two separate entities.
@@
 Either way, this work implies that the corona and the outflow are connected through some physical process.
-Considering our findings and following the models of Markoff et al.
-(2001b); Markoff & Nowak (2004), we argue that the corona and the jet acceleration region are coupled, with continuous feedback between them (Malzac, 2014).
+Considering our findings and following the models of Markoff2001; Markoff2004, we argue that the corona and the jet acceleration region are coupled, with continuous feedback between them (Malzac2014).
 A magnetic field likely originating from the accretion disk, or threaded in the event horizon could be the mechanism behind this coupling.
-Parts of the corona could be ejected out far beyond the lamp-post height assumptions and possibly interact with the first acceleration region or jet base through the magnetic field lines (Narayan et al., 2003; Igumenshchev, 2008; Tchekhovskoy et al., 2011; Janiuk & James, 2022).
-An alternative explanation could be that a change in the accretion rate close to the HS to SS transition (point of transient jet eject) makes fast-moving matter clump into slow-moving plasma, creating a shock in the accretion disk (Fender et al., 2004c; Narayan et al., 2012; Janiuk & James, 2022).
-This shock changes the accompanying magnetic field configuration (Narayan et al., 2003; Igumenshchev, 2008; Tchekhovskoy et al., 2011; Janiuk & James, 2022), and leads to a change in the corona geometry, possibly from compact to the lamp-post structure, which leads to a change in the coronal height.
+Parts of the corona could be ejected out far beyond the lamp-post height assumptions and possibly interact with the first acceleration region or jet base through the magnetic field lines (Narayan2003; Igumenshchev2008; Tchekhovskoy2011; Janiuk2022).
+An alternative explanation could be that a change in the accretion rate close to the LHS to HSS transition (point of transient jet eject) makes fast-moving matter clump into slow-moving plasma, creating a shock in the accretion disk (Fender2004a; Narayan2012).
+This shock may change the accompanying magnetic field configuration and lead to a change in the corona geometry, possibly from compact to the lamp-post structure, which leads to a change in the coronal height.
 This also suggests that the corona might have a more complex geometry, and it could be a combination of both an extended corona and the near-point source.
-Rapid changes to the coronal geometry as the jet transitions to transient have already been observed in Swift J1727.8–1613 (Podgorný et al., 2024).
-The authors also reported a return to pre-outburst polarization levels upon re-entering the HS, indicating that the coronal geometry and nature were similar to those before the outburst (Podgorný et al., 2024).
-Furthermore, a drastic decrease in polarization during the SS has been widely observed (Mastroserio et al., 2024; Svoboda et al., 2024).
-The polarization findings also lend support to the notion of a non-trivial coronal geometry; a radially extended portion is especially indicated by polarization (Veledina et al., 2023; Krawczynski et al., 2022; Ingram et al., 2024; Steiner et al., 2024).
+Rapid changes to the coronal geometry as the jet becomes transient have already been observed in Swift J1727.8–1613 (Podgorny2024).
+The authors also reported a return to pre-outburst polarization levels upon re-entering the LHS, indicating that the coronal geometry and nature were similar to those before the outburst (Podgorny2024).
+Furthermore, a drastic decrease in polarization during the HSS has been observed (Mastroserio2024; Soboda2024).
+The polarization findings also lend support to the notion of a non-trivial coronal geometry; a radially extended portion is especially indicated by polarization (Veledina_2023_ixpe; cyg_ixpe; Ingram_2024; Steiner2024) However, it should be noted that interpretation of the polarization measurements hinges on simulation-based models that only consider certain geometries.
+Although some reverberation studies study suggest a small corona, others have also suggested the possibility of a corona that is radially extended along the disk (ma2023; alabarta2025).
+Reconciling these findings with spectral modeling results suggesting a more compact or vertically extended coronal geometry is challenging.
+It has been suggested (e.g.
+alabarta2025) that the corona both begins and ends an outburst with a compact component and a large, horizontally extended component, while the large component takes on a jet-like shape at the peak of the outburst.
 Our analysis cannot determine whether the jet ejection or the coronal height increase occurred first or whether the corona is entwined with the compact jet or the transient jet.
-Jet ejection times are difficult to pinpoint (Wood et al., 2021; Russell et al., 2020; Carotenuto et al., 2021).
+Jet ejection times are difficult to pinpoint (Wood_2021; Russell_2020; car).
 Additionally, our X-ray spectral epochs are too irregular to track the exact onset of coronal height changes, making it challenging to identify the cause of these geometric changes.
@@
 Acknowledgements
+We sincerely thank the referee for their thorough and constructive review.
+We appreciate the time and effort they put into evaluating our manuscript, and we have carefully considered all comments and suggestions.
 This work is partly supported by the NSF REU Program (Award Number 2244344) at Montana State University.
@@
 JC and RS acknowledge support from the Leverhulme Trust grant RPG-2023-240.
+| MAXI J1348–630 | ||||||
+| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp.
+Data 333Ni=NICER, S = Swift/XRT | Supp.
+ObsID | Corona Heights ()444p means model pegged height at minimum value,-p means lower limit is minimum value |
+| 1 | 58515.13 | LHS | 80402315002 | - | ||
+| 2 | 58515.60 | LHS | 80402315004 | - | ||
+| 3 | 58521.07 | HIMS | 80402315006 | S | 00011107004 | |
+| 4 | 58525.26 | SIMS | 80402315008 | - | ||
+| 5 | 58550.64 | HSS | 80402315010 | S | 00088843002 | |
+| 6 | 58577.03 | HSS | 80402315012 | Ni | 2200530118 | |
+| 7 | 58655.71 | LHS | 80502304002 | Ni | 2200530161 | |
+| 8 | 58660.50 | LHS | 80502304004 | - | ||
+| 9 | 58672.81 | LHS | 80502304006 | Ni | 2200530169 | |
+| : | 1.07 | |||||
+| MAXI J1820+070 Group 1 | ||||||
+| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
+| 1 | 58191.94 | LHS | 90401309002 | - | ||
+| 2 | 58198.04 | LHS | 90401309004 | - | ||
+| 3 | 58198.39 | LHS | 90401309006 | Ni | 1200120106 | |
+| 4 | 58201.53 | LHS | 90401309008 | Ni | 1200120109 | |
+| 5 | 58201.86 | LHS | 90401309010 | Ni | 1200120109 | |
+| 6 | 58212.22 | LHS | 90401309012 | Ni | 1200120120 | |
+| 7 | 58224.96 | LHS | 90401309013 | Ni | 1200120130 | |
+| 8 | 58225.30 | LHS | 90401309014 | Ni | 1200120131 | |
+| : | 1.01 | |||||
+| MAXI J1820+070 Group 2 | ||||||
+| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
+| 9 | 58242.27 | LHS | 90401309016 | Ni | 1200120144 | |
+| 10 | 58255.15 | LHS | 90401309018 | - | ||
+| 11 | 58255.91 | LHS | 90401309019 | S | 00088657004 | |
+| 12 | 58259.33 | LHS | 90401324002 | Ni | 1200120155 | |
+| 13 | 58297.20 | LHS | 90401309021 | Ni | 1200120189 | |
+| 14 | 58306.38 | SIMS | 90401309023 | Ni | 1200120198 | |
+| 15 | 58315.02 | HSS | 90401309025 | Ni | 1200120208 | |
+| 16 | 58327.04 | HSS | 90401309026 | - | ||
+| 17 | 58327.05 | HSS | 90401309027 | - | ||
+| : | 0.98 | |||||
+| MAXI J1820+070 Group 3 | ||||||
+| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
+| 18 | 58343.61 | HSS | 90401309029 | - | ||
+| 19 | 58349.82 | HSS | 90401309031 | Ni | 1200120240 | |
+| 20 | 58389.35 | Int | 90401309033 | Ni | 1200120267 | |
+| 21 | 58397.30 | LHS | 90401309035 | - | ||
+| 22 | 58404.95 | LHS | 90401309037 | - | ||
+| 23 | 58420.07 | LHS | 90401309039 | Ni | 1200120290 | |
+| 0.99 | ||||||
+| MAXI J1535–571 Group 1 | ||||||
+| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
+| 1 | 58004.35 | LHS | 90401309029 | S | 00010264003 | |
+| 2 | 58008.72 | HIMS | 80302309002 | Ni | 1050360104 | |
+| 3 | 58010.26 | HIMS | 80402302002 | Ni | 1050360106 | |
+| 4 | 58010.57 | HIMS | 80402302004 | - | ||
+| 5 | 58013.22 | HIMS | 80302309004 | Ni | 1050360109 | |
+| 6 | 58013.48 | HIMS | 80302309006 | Ni | 1050360109 | |
+| 7 | 58016.87 | SIMS | 80302309010 | - | ||
+| 8 | 58017.21 | SIMS | 80302309010 | - | ||
+| : | 1.07 | |||||
+| MAXI J1535–571 Group 2 | ||||||
+| Epoch | Date (MJD) | State | NuSTAR ObsID | Supp. Data | Supp. ObsID | Corona Heights () |
+| 9 | 58023.11 | SIMS | 80402302006 | - | ||
+| 10 | 58032.71 | SIMS | 80302309012 | - | ||
+| 11 | 58034.26 | Int | 80402302008 | - | ||
+| 12 | 58048.99 | LHS | 80302309014 | S | 00088245004 | |
+| 13 | 58050.98 | LHS | 80402302009 | S | 00088246001 | |
+| 14 | 58051.38 | LHS | 80402302010 | - | ||
+| 15 | 58090.43 | HSS | 90301327002 | - | ||
+| 16 | 58090.77 | HSS | 90301327004 | - | ||
+| : | 1.15 | |||||
+=-1in
+| MAXI J1348–630 | ||||||||||
+| Model | Parameter | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 | Epoch 9 |
+| tbabs | nH ( atoms cm-2)555nH, a, i, and AFe tied across all epochs in each group | |||||||||
+| diskbb | Tin (keV) | |||||||||
+| Nd | ||||||||||
+| relxilllp | h (rg) 666-p means lower bound is lower limit,+p means upper bound at upper limit,p means parameter pegged at limit | |||||||||
+| a | ||||||||||
+| i (degrees) | ||||||||||
+| Rin (rg) | ||||||||||
+| log | ||||||||||
+| AFe | ||||||||||
+| Ecut (keV) | ||||||||||
+| Nr | ||||||||||
+| Constant | FNuB 777Multiplicative factor for NuSTAR FPMA | |||||||||
+| FSup888Multiplicative factor for supplemental data, if used. | ||||||||||
+| = 1.07 | ||||||||||
+=-1in
+| MAXI J1820+070 Group 1 | |||||||||||
+| Model | Par | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 | ||
+| tbabs | nH ( atoms cm-2) | ||||||||||
+| diskbb | Tin (keV) | ||||||||||
+| Nd | |||||||||||
+| bbody | kT (keV) | ||||||||||
+| Nbb | |||||||||||
+| relxilllp | h (rg) | ||||||||||
+| a | |||||||||||
+| i (degrees) | |||||||||||
+| Rin (rg) | |||||||||||
+| log | |||||||||||
+| AFe | |||||||||||
+| Ecut (keV) | |||||||||||
+| Nr | |||||||||||
+| Constant | FNuB | ||||||||||
+| FSup | |||||||||||
+| = 1.01 | |||||||||||
+=-1in
+| MAXI J1820+070 Group 1 with Hydrogen Column Density Fixed | |||||||||||
+| Model | Par | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 | ||
+| diskbb | nH ( atoms cm-2) | ||||||||||
+| Tin (keV) | |||||||||||
+| Nd | |||||||||||
+| bbody | kT (keV) | ||||||||||
+| Nbb | |||||||||||
+| relxilllp | h (rg) | ||||||||||
+| a | |||||||||||
+| i (degrees) | |||||||||||
+| Rin (rg) | |||||||||||
+| log | |||||||||||
+| AFe | |||||||||||
+| Ecut (keV) | |||||||||||
+| Nr | |||||||||||
+| Constant | FNuB | ||||||||||
+| FSup | |||||||||||
+| = 1.06 | |||||||||||
+=-1.0in
+| MAXI J1820+070 Group 2 | ||||||||||
+| Model | Par | Epoch 9 | Epoch 10 555Disk not detected | Epoch 11 666Disk and blackbody component not detected | Epoch 12 | Epoch 13 | Epoch 14 | Epoch 15 | Epoch 16 | Epoch 17 |
+| tbabs | nH ( atoms cm-2) | |||||||||
+| diskbb | Tin (keV) | |||||||||
+| Nd () | ||||||||||
+| bbody | kT (keV) | |||||||||
+| Nbb | ||||||||||
+| relxilllp | h (rg) | |||||||||
+| a | ||||||||||
+| i (degrees) | ||||||||||
+| Rin (rg) | ||||||||||
+| log | ||||||||||
+| AFe | ||||||||||
+| Ecut (keV) | ||||||||||
+| Nr | ||||||||||
+| Constant | FNuB | |||||||||
+| FSup | ||||||||||
+| = 0.98 | ||||||||||
+| MAXI J1820+070 Group 3 | |||||||
+| Model | Par | Epoch 18 | Epoch 19 | Epoch 20 | Epoch 21 | Epoch 22 | Epoch 23 |
+| tbabs | nH ( atoms cm-2) | ||||||
+| diskbb | Tin (keV) | ||||||
+| Nd | |||||||
+| bbody | kT (keV) | ||||||
+| Nbb | |||||||
+| relxilllp | h (rg) | ||||||
+| a | |||||||
+| i (degrees) | |||||||
+| Rin (rg) | |||||||
+| log | |||||||
+| AFe | |||||||
+| Ecut (keV) | |||||||
+| Nr | |||||||
+| Constant | FNuB | ||||||
+| FSup | |||||||
+| = 0.99 | |||||||
+| MAXI J1535–571 Group 1 | |||||||||
+| Model | Par | Epoch 1 | Epoch 2 | Epoch 3 | Epoch 4 | Epoch 5 | Epoch 6 | Epoch 7 | Epoch 8 |
+| tbabs | nH ( atoms cm-2) | ||||||||
+| diskbbb | Tin (keV) | ||||||||
+| Nd | |||||||||
+| relxilllp | h (rg) | ||||||||
+| a | |||||||||
+| i (degrees) | |||||||||
+| Rin (rg) | |||||||||
+| log | |||||||||
+| AFe | |||||||||
+| Ecut (keV) | |||||||||
+| Nr | |||||||||
+| Constant | FNuB | ||||||||
+| FSup | |||||||||
+| = 1.07 | |||||||||
+| MAXI J1535–571 Group 2 | |||||||||
+| Model | Par | Epoch 9 | Epoch 10 | Epoch 11 | Epoch 12 | Epoch 13 | Epoch 14 | Epoch 15 | Epoch 16 |
+| tbabs | nH ( atoms cm-2) | ||||||||
+| diskbb | Tin (keV) | ||||||||
+| Nd | |||||||||
+| relxilllp | h (rg) | ||||||||
+| a | |||||||||
+| i (degrees) | |||||||||
+| Rin (rg) | |||||||||
+| log | |||||||||
+| AFe | |||||||||
+| Ecut (keV) | |||||||||
+| Nr | |||||||||
+| Constant | FNuB | ||||||||
+| FSup | |||||||||
+| = 1.15 | |||||||||